40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 15 15:11:38 2011, kiwamu, Update, SUS, f2a filters on BS and PRM PRMf2a.pngBS_f2a.png
    Reply  Fri Sep 16 16:29:05 2011, kiwamu, Update, SUS, f2a filters on SRM F2ASRM_Sep16.png
Message ID: 5435     Entry time: Fri Sep 16 16:29:05 2011     In reply to: 5417
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: f2a filters on SRM 

New f2a filters were installed on SRM.

The lock of DRMI should be more stable than last night.

Quote from #5417

Once the SRM oplev project settles down, I will adjust the f2a filters on SRM too.

 

Attachment 1: F2ASRM_Sep16.png  42 kB  | Hide | Hide all
F2ASRM_Sep16.png
ELOG V3.1.3-